# Anomalous Criticality of Absorbing State Transition toward Jamming

### v1

@@ Abstract
Abstract: The jamming transition is traditionally regarded as a geometric transition governed by static contact networks. Recently, dynamic phase transitions of athermal particles under periodic shear provide a new lens on this problem, leading to a conjecture that jamming transition corresponds to an absorbing-state transition within the Manna (conserved directed percolation) universality class. Here, by re-examing the biased random organization model, a minimal model for particles under periodic shearing that the conjecture bases on, we uncover several criticality anomalies at high density beyond the description of Manna universality class. In three-dimensional monodisperse systems, crystallization disrupts the absorbing transition, while in dense binary mixtures, a distinct transition from absorbing to active-glass states emerges, signifying a new universality class of dynamic phase transitions. Closer to jamming point, the quenched heterogeneity in the contact network smears the dynamic transition via Griffiths effects and drives the system toward heterogeneous directed percolation. We further proposal a field theory with fractional time dynamics that unifies these phenomena, establishing a theoretical framework linking jamming, disorder, and dynamic criticality.
I Significance
The jamming transition, when materials like sand or foam suddenly become rigid, is traditionally viewed as a geometric packing problem. Recent ideas suggest it’s also tied to a criticality of dynamic processes, like how crowded particles stop moving under repeated shaking or shearing. This work puts these ideas to the test and reveal a more complex and fascinating story: at intermediate high density, a new universality class emerges, marking a transition from absorbing to active-glass states, while near the jamming point, heterogeneity induces Griffiths effects smearing out the dynamic critical point. Our work reveals complicated interplay between static and dynamic phase transitions and sheds lights on the dynamic criticality of high-dimensional disordered systems, like learning processes of artificial neuron networks.
@@ II Introduction
The jamming transition, a transition from a mechanically soft state to a rigid state, widely exists in soft matter systems [1, 2, 3, 4]. The static packing of spheres provides a natural route to study the jamming transition, in which the random close packing (RCP) state is introduced to describe the most compact state that spheres can be arranged in a random way [5, 6, 7]. Although the RCP state can be protocol-dependent [8, 9], it is still widely believed to be a signature of the jamming transition [10, 11, 12, 13, 14, 15, 16].
@@
Another path to tackle the jamming problem is from the dynamic side. It was first found that when dilute suspended colloidal particles are subjected to periodic shearing, a reversible to irreversible transition, or an absorbing phase transition, can occur upon increasing the shearing amplitude [17, 18, 19]. As system density increases, the critical shearing amplitude becomes smaller, making the reversible to irreversible transition coincide with the jamming transition [20, 19, 21, 22, 23, 24, 25, 26, 4, 27, 28]. Nevertheless, directly studying the criticality of such dynamic transition at high density is challenging, due to the diverging relaxation time and long-range correlation between particles [23, 22, 19, 21]. Theoretically, the many-body dynamics of particles under periodic shearing can be well captured by a minimal model, i.e., the random organization (RO) model [18]. Recent works have demonstrated that the biased random organization (BRO) model, a more realistic version of the RO model, can generate the RCP state at the critical point of the absorbing transition when the displacement size (corresponding to shearing amplitude) is vanishing $\varepsilon\rightarrow 0$ in dimensions $d=3$ to $5$ [29, 30, 31]. The RO/BRO model in the dilute regime has been proven to belong to the Manna universality class, or conserved directed percolation [32, 33, 34]. In Ref. [29, 30, 35], it was suggested that the BRO model also resides in the Manna universality class in the close packing limit ($\varepsilon\rightarrow 0$). Nevertheless, a comprehensive and detailed study of the dynamic criticality of the BRO model in the high density regime is needed to fully resolve this issue.
@@
Figure 1: (a) Schematic of the BRO model: overlapping particles (red) become active and take pairwise repulsive displacements. (b) $f_{a}^{\infty}$ as a function of packing fraction $\varphi$ at $\epsilon=2.0\sigma$. (c) Evolution of $f_{b}(t)$ for systems at $\varepsilon=2.0$ under various $\varphi$. The dashed line represents the power law $t^{-0.45}$. Here $N=65536$. (d) Critical packing fractions $\varphi_{c}$ for 2D binary (red line), 3D binary (orange dots), 3D monodisperse (blue circles) and 4D monodisperse (green open squares) systems at different $\varepsilon$. Asterisks mark the extrapolated $\varphi_{c}$ for $\varepsilon\rightarrow 0$, i.e., 0.840 (2D) and 0.638 (3D) and 0.455 (4D), which is consistent with the RCP.
@@
In this work, we show that the BRO model exhibits anomalous critical behaviors at high density, which deviates from Manna universality class. In 3D monodisperse sphere systems, we find that the absorbing transition is interrupted by crystallization. In binary mixture systems without crystallization, an unreported universality of dynamic phase transition, i.e., absorbing to active-glass transition, is identified, which is characterized by critical exponents different from the Manna universality. With further increasing the critical density by decreasing $\varepsilon$, dynamic heterogeneity and Griffiths effects emerge, which smooth the dynamic transition and broaden the critical point into a pseudo-critical regime. As $\varepsilon\rightarrow 0$, the dynamic transition becomes equivalent to heterogeneous directed percolation (DP) due to particle immobility set by the jamming. These results indicates that while the BRO model can generate RCP configurations, the criticality of the transition is fundamentally reshaped by the RCP state. This discourages a direct mapping of the jamming transition to an absorbing phase transitions. At last, we also propose a field theory with fractional time dynamics which can describe the transition from absorbing to active-glass transition and heterogeneous DP in the same framework, revealing a complicated interplay between dynamic criticality and spatial annealed/quenched disorder. Our findings can be a starting point to study the criticality of dynamic phase transition in more complicate disordered systems, like the learning process of artificial neuron networks [31, 36].
@@ III Model and simulations
In the BRO model, particles are first randomly distributed in the space of dimension $d$. In each discrete time step, if a particle is isolated without overlapping with other particles, it is considered inactive and remains stationary. On the contrary, if two particles overlap, they are considered as active and a pair of repulsive displacements with opposite direction are assigned to each other. The displacement pair has the same magnitude randomly chosen from uniform range of $0,\varepsilon/2]$, as shown in Fig. 1(a). If a particle overlaps multiple particles, pairwise displacements are superimposed [29, 30, 31]. We simulate $N$ spherical particles in a square box of length $L$ with periodic boundary conditions at packing fraction $\varphi$. Both the monodisperse particle system and equal-number binary mixture of small and large particles are considered. In the latter, the particle diameter ratio is $\sigma_{M}/\sigma_{m}=1:\sqrt{2}$ for 2D systems. For 3D systems, a larger particle diameter ratio $\sigma_{M}/\sigma_{m}=1:1.6$ is also considered. The small particle diameter in each system is chosen as the unit length $\sigma$.
@@ IV Results
Figure 3: (a, e) Evolution of $f_{b}(t)$ at different $\varphi$, where Griffiths regime is painted in gray. (b, f) Evolution of $f_{b}(t)$ at different $N$. And $\varphi=0.8128$(upper) and $0.4920$ (lower). (c, g) Relationship between index $\kappa$ and $\varphi$. (d, h) Distributions of normalized activity duration $f_{ad}$ at $\varphi=0.7838,\varepsilon=0.2\sigma$(upper) and $0.4980,\varepsilon=1.5\sigma$, $\Delta\epsilon=1.2\sigma$, $l=1\sigma$ (lower). (a, b) are for the original BRO model with $\varepsilon=0.1\sigma$. (e, f) are for modified BRO models with quenched-disorder at $\varepsilon=1.5\sigma$, $\Delta\epsilon=1.2\sigma$ and $l=2\sigma$. For all systems except (b,f), $N=8192$.
Nevertheless, for high density 3D monodispersed systems with $\varepsilon<0.8\sigma$, we find that the absorbing transition is interfered by crystallization, thus does not exhibit standard critical phenomena (see Fig. S1). Although binary systems can avoid crystallization, their critical behaviors do not match the Manna universality. In Fig. 2(a, b, d, e), we show the finite-size scaling of $f_{a}(t)$, $P_{s}(t)$, $f_{a}^{\infty}(\tau)$ and $t^{*}(\tau)$ for 2D binary systems with $\varepsilon=0.8\sigma$ at $\varphi_{c}=0.6324$. This finite-size scaling is different from the typical BRO model with a large step size ($\varepsilon=1.5\sigma$) (see Fig. S3). The obtained critical exponents $\beta$, $\nu_{\parallel}$, $\alpha$, $z$ as functions of $\varepsilon$ are shown in Fig. 2(c, f), where for $\epsilon>1.0\sigma$ (Regime I), these critical exponents are consistent with the Manna universality. When $0.3\sigma<\varepsilon<1.0\sigma$ (Regime II), we find non-monotonic behaviors in $\nu_{\perp}$, $\alpha$, and $z$ inconsistent with the Manna universality. Further decreasing $\varepsilon$ below $0.3\sigma$ (Regime III), the failure of standard finite size scaling and the uncertainty of the critical point are observed. It should be noted that the exponent $\beta$ does not change significantly in Regimes I and II. The measured $z$, $z^{\prime}$, and $z^{*}$ are consistent, and $\nu_{\perp}=z/\nu_{\parallel}$ also changes with $\varepsilon$ (see Fig. S2). In Fig. S2-6, we also present the exponents measured in the 3D system under two different particle diameter ratios, as well as the mono-dispersed 4D system. We find that the general behavior of the critical exponent and finite size scaling remain unchanged. This indicates that the anomalous criticality of the BRO model are robust at high density, insensitive to the dimensionality and particle diameter ratio.
@@
We first focus on the Regime I where the step size $\varepsilon$ is larger than the particle size. In this regime, the motion of active particles is not constrained by neighbouring particles. Thus, the mean-square-displacement (MSD) of active particles shows diffusion scaling as shown in Fig. 2(g). Here, the solid and open symbols represent systems above and below the critical point, respectively. The diffusion scaling in sub-critical systems indicates that the absorbing transition is essentially a non-local searching process for a non-overlapping configurations, consistent with Manna (CDP) universality [37, 38].
As $\varepsilon$ decreases, $\varphi_{c}$ increases. Due to the repulsive displacement in the BRO model, the particles are caged by their neighbours. This effect is shown in Fig. 2(h) for systems with $\varepsilon=0.5\sigma$, where one can find a early-time plateau or sub-diffusion regime in MSD for system near the critical point, a characteristic of the glass-forming liquids [39, 40, 41, 42]. In Fig. 2(i), we plot the measured diffusion coefficient $D$ of active particle and $f_{a}^{\infty}$ as a function of $\varphi$ under different $\varepsilon$. We can clearly distinguish active states with $D$ as low as $10^{-6}\sigma^{2}/\tau_{0}$ for $\varepsilon\lesssim\sigma$. The decoupling between $f_{a}^{\infty}$ and $D$ suggests that in this regime (Regime II), the absorbing phase transition is essentially an absorbing to active-glass transition, where particles can only find a non-overlapping absorbing configuration through local arrangements instead of non-local searching.
The fundamental difference between these two scenarios can be distinguished by field theory. For Manna (CDP) universality, it can be described by an non-conserved field $\rho({\bf{r}},t)$ (active particle density field) coupling with a non-diffusive conserved field $\psi({\bf{r}},t)$ (total density field)[43, 44], namely,
@@
with $\theta=1$ indicating the evolution of $\psi(t)$ field is a result of diffusion of $\rho(t)$ field. Note that $a$, $b$, $w$ are fixed parameters determining the critical point[44]. The last term in Eq.(4) is the multiplicative noise term accounting for absorbing phase transition. In the active-glass state, however, the dynamic of $\psi(t)$ field in Eq.(4) is sub-diffusive, corresponding to fractional time derivative with $0<\theta<1$[45, 46]. Obviously, the criticality with fractional time dynamics distinguish itself from Manna universality. The observed continuous variation of critical exponents in this regime might be due to the variation of $\theta$. Nevertheless, numerical solution and renormalization-group analysis of this fractional time equation with multiplicative noise is currently challenging [45, 46].
@@
When $\varepsilon$ is further reduced below $0.3\sigma$ (Regime III), the critical density of the system would approach to that of RCP state. In this regime, caging-breaking events becomes unlikely and the topology of particle network remains unchanged in the activation process. We refer this state as the pre-jamming state. Different from the conventional absorbing phase transition, we find $f_{b}(t)$ in this state relaxes according to a continuous power law
@@
In the Griffiths phase, the quenched disorder generates strong dynamic heterogeneity in the relaxation process. This can be visualized by the spatial distribution of normalized activity duration $f_{ad}$, i.e., the fraction of time that a particle stays in an active state relative to the active duration of the whole system, starting from a perturbed absorbing configuration. In Fig. 3(d), we show these distributions for systems with $\varepsilon=0.2\sigma$. One can observe non-uniform network structures of activity, resembling the dynamic heterogeneity of a supercooled liquids or the force chain of the jammed state [53, 54, 55]. In Regime II, one can also find relative weak dynamic heterogeneity (Fig. S9), which would be smoothed out by further increasing $\varepsilon$ to Regime I.
In the Griffiths phase, the finite-size effect of $f_{b}(t)$ and $P_{s}(t)$ near criticality are also reversed, i.e., at the same density near the transition, small systems stay longer in the active state than large systems. This is best shown in Fig. 3(b) and Fig. 4(a), as well as in Fig. S3. This abnormality is nevertheless consistent with the fact that small systems are more likely be jammed near the jamming point [56, 57]. This suggests that close-packing significantly influences the criticality of absorbing phase transition in this regime. Further support comes from the structure factor $S(q)$ for systems with $\epsilon\rightarrow 0$, which does not follow the $q^{0.45}$ scaling of the Manna model (see Fig. S8).
To further prove that the above phenomena are results of Griffiths effect, we manually add quenched disorder in systems located at Regime I, by making $\varepsilon$ spatially dependent. More explicitly, the system is divided into cells of size $l$. Each cell $(i,j)$ has its own $\varepsilon_{i,j}$ set as
@@
Figure 4: (a, c) Finite size effects of survival probability of dynamic trajectory $P_{s}(t)$. (d, f) Distributions of normalized activity duration. (a, b) are for the original BRO models near the transition point at $\varepsilon=0.05\sigma$ and $\varphi=0.8380$. (c, d) are for the heterogeneous contact model at $P_{e}=0.795$ and $k=1$ based on the same configuration as in (b).
@@
Finally, we study the critical behaviors of the BRO model in the close packing limit ($\varepsilon\rightarrow 0$). In this limit, the particle configuration can be regarded as static during the relaxation process. From the point of field theory, this scenario corresponds to $\theta=0$ in Eq.(4), where the evolution of density field $\psi$ is completely forbidden. Thus, the density field $\psi$ acts as pure quenched disorder in the dynamics of $\rho$ field, essentially changing the CDP universality to the directed percolation (DP) universality [50, 49, 47]. The classical contact model is the simplest realization of DP [59, 60]. To test this idea, we simulate a contact model on the networks composed of nearly jammed particles generated by the BRO model at $\varepsilon=0.05\sigma$, where each particle represents a node. We find that the topology of the contact network alone does not affect the critical behaviours of systems. However, if the infection rate of a node is proportional to the activity durations of the corresponding particle (see the Method section for details), the activity distribution of this contact model exhibits dynamic heterogeneity resembling that of the high density BRO model, as shown in Fig. 4(b, d). Correspondingly, the finite size scaling of $P_{s}(t)$ in Fig. 4(c) also exhibits similar anomalous behaviors as the BRO model in the close packing limit Fig. 4(a). This similarity proves that the critical behaviors of the BRO model in the nearly jammed configuration corresponds to DP on hierarchic heterogeneous network. Moreover, the Griffith effect becomes stronger as the heterogeneity increases in DP model (see Fig. S11). It should be noted that our previous modified BRO model with cell-size quenched-disorder fails to account Griffith effect at the closed packing limit, because the correlation in this point span the system size, leaving the topology of heterogeneous network play the dominant role.

### v2

@@ Abstract
Abstract: The jamming transition is traditionally regarded as a geometric transition governed by static contact networks. Recently, dynamic phase transitions of athermal particles under periodic shear provide a new lens on this problem, leading to a conjecture that jamming transition corresponds to an absorbing-state transition within the Manna (conserved directed percolation) universality class. Here, by re-examining the biased random organization model, a minimal model for particles under periodic shearing that the conjecture is based on, we uncover several criticality anomalies at high density at odds with Manna universality class. In three-dimensional monodisperse systems, we find crystallization disrupts the absorbing transition, while in dense binary mixtures, a distinct transition from absorbing to active-glass states emerges, signifying a new universality class of dynamic phase transition. Closer to the jamming point, the quenched heterogeneity in the contact network smears the dynamic transition via Griffiths effects and drives the system toward heterogeneous directed percolation. We propose a field theory with fractional time dynamics that unifies these phenomena, establishing a theoretical framework linking jamming, disorder, and dynamic criticality.
I Popular Summary
The jamming transition, when materials like sand or foam suddenly become rigid, is traditionally viewed as a geometric packing problem. Recent ideas suggest it’s also tied to a criticality of dynamic processes, like how crowded particles stop moving under repeated shaking or shearing. This work puts these ideas to the test and reveals a more complex and fascinating story: at intermediate high density, a new universality class emerges, marking a transition from absorbing to active-glass states, while near the jamming point, heterogeneity induces Griffiths effects smearing out the dynamic critical point. Our work reveals complicated interplay between static and dynamic phase transitions and sheds light on the dynamic criticality of high-dimensional disordered systems, like learning processes of artificial neural networks.
@@ II Introduction
The jamming transition, a transition from a mechanically soft state to a rigid state, widely exists in soft matter systems [1, 2, 3, 4]. The static packing of spheres provides a natural route to study the jamming transition, in which the random close packing (RCP) state is introduced to describe the most compact state that spheres can be arranged in a random way [5, 6, 7]. Although the RCP state can be protocol-dependent [8, 9], it is still widely believed to be a signature of the jamming transition [10, 11, 12, 13, 14, 15, 16, 17].
@@
Another path to tackle the jamming problem is from the dynamic side. It was first found that when dilute suspended colloidal particles are subjected to periodic shearing, a reversible to irreversible transition, or an absorbing phase transition, can occur upon increasing the shearing amplitude [18, 19, 20]. As system density increases, the critical shearing amplitude becomes smaller, making the reversible to irreversible transition coincide with the jamming transition [21, 37, 20, 22, 23, 24, 25, 26, 27, 4, 28, 29]. Nevertheless, directly studying the criticality of such a dynamic transition at high density is challenging, due to the divergent relaxation time and long-range correlation near the jamming point [24, 23, 20, 22]. Theoretically, the many-body dynamics of particles under periodic shearing can be well captured by a minimal model, i.e., the random organization (RO) model [19]. Recent works have demonstrated that the biased random organization (BRO) model, a more realistic version of the RO model, can generate the RCP state at the critical point of the absorbing transition when the displacement size (corresponding to shearing amplitude) vanishes $\varepsilon\rightarrow 0$ in dimensions $d=3$ to $5$ [30, 31, 32]. The RO/BRO model in the dilute regime has been proven to belong to the Manna universality class, or conserved directed percolation (CDP) [33, 34, 35, 36, 37]. In Ref. [30, 31, 38], it was suggested that the BRO model also resides in the Manna universality class in the close packing limit ($\varepsilon\rightarrow 0$). Nevertheless, a comprehensive and detailed study of the dynamic criticality of the BRO model in the high density regime is needed to fully resolve this issue.
@@
Figure 1: (a) Schematic of the BRO model: overlapping particles (red) become active and take pairwise repulsive displacements. (b) $f_{a}^{\infty}$ as a function of packing fraction $\varphi$ at $\varepsilon=2.0~\sigma$. (c) Evolution of $f_{b}(t)$ for systems at $\varepsilon=2.0~\sigma$ under various $\varphi$. The dashed line represents the power law $t^{-0.45}$. Here $N=65536$. (d) Critical packing fractions $\varphi_{c}$ for 2D binary (red line), 3D binary (orange dots), 3D monodisperse (blue circles) and 4D monodisperse (green open squares) systems at different $\varepsilon$. Asterisks mark the extrapolated $\varphi_{c}$ for $\varepsilon\rightarrow 0$, i.e., 0.840 (2D) and 0.638 (3D) and 0.455 (4D), which is consistent with the RCP.
@@
In this work, we show that the BRO model exhibits anomalous critical behaviors at high density, which deviates from Manna universality class. In 3D monodisperse sphere systems, we find that the absorbing transition is interrupted by crystallization. In binary mixture systems without crystallization, an unreported universality of dynamic phase transition, i.e., absorbing to active-glass transition, is identified, which is characterized by critical exponents different from the Manna universality. With further increasing the critical density by decreasing $\varepsilon$, dynamic heterogeneity and Griffiths effects emerge, which smooth the dynamic transition and broaden the critical point into a pseudo-critical regime. As $\varepsilon\rightarrow 0$, the dynamic transition becomes equivalent to heterogeneous directed percolation (DP) due to particle immobility set by the jamming. These results indicates that while the BRO model can generate RCP configurations, the criticality of the transition is fundamentally reshaped by the RCP state. This discourages a direct mapping of the jamming transition to an absorbing phase transitions. At last, we also propose a field theory with fractional time dynamics which can describe the transition from absorbing to active-glass transition and heterogeneous DP in the same framework, revealing a complicated interplay between dynamic criticality and spatial annealed/quenched disorder. Our findings can be a starting point to study the dynamic criticality of more complicate disordered systems, like the learning process of artificial neural networks [32, 39].
@@ III Model and simulations
In the BRO model, particles are first randomly distributed in the space of dimension $d$. In each discrete time step, if a particle is isolated without overlapping with other particles, it is considered inactive and remains stationary. On the contrary, if two particles overlap, they are considered as active and a pair of repulsive displacements with opposite direction are assigned to each other. The displacement pair has the same magnitude randomly chosen from uniform range of $0,\varepsilon/2]$, as shown in Fig. 1(a). If a particle overlaps multiple particles, pairwise displacements are superimposed [30, 31, 32]. We simulate $N$ spherical particles in a hypercubic box of length $L$ with periodic boundary conditions at packing fraction $\varphi$. Both the monodisperse particle system and equal-number binary mixture of small and large particles are considered. In the latter, the particle diameter ratio is $\sigma_{M}/\sigma_{m}=1:\sqrt{2}$ for 2D systems. For 3D systems, a larger particle diameter ratio $\sigma_{M}/\sigma_{m}=1:1.6$ is also considered. The small particle diameter in each system is chosen as the unit length $\sigma$.
@@ IV Results
Figure 3: (a, e) Evolution of $f_{b}(t)$ at different $\varphi$, where Griffiths regime is painted in gray. (b, f) Evolution of $f_{b}(t)$ at different $N$. And $\varphi=0.8128$(upper) and $0.4920$ (lower). (c, g) Relationship between index $\kappa$ and $\varphi$. (d, h) Distributions of normalized activity duration $f_{ad}$ at $\varphi=0.7838,\varepsilon=0.2\sigma$(upper) and $0.4980,\varepsilon=1.5\sigma$, $\Delta\varepsilon=1.2\sigma$, $l=1\sigma$ (lower). (a, b) are for the original BRO model with $\varepsilon=0.1\sigma$. (e, f) are for modified BRO models with quenched-disorder at $\varepsilon=1.5\sigma$, $\Delta\varepsilon=1.2\sigma$ and $l=2\sigma$. For all systems except (b,f), $N=8192$.
Nevertheless, for high density 3D monodispersed systems with $\varepsilon<0.8\sigma$, we find that the absorbing transition is disrupted by crystallization, thus does not exhibit standard critical phenomena (see Fig. S1). Although binary systems can avoid crystallization, their critical behaviors do not match the Manna universality. In Fig. 2(a, b, d, e), we show the finite-size scaling of $f_{a}(t)$, $P_{s}(t)$, $f_{a}^{\infty}(\tau)$ and $t^{*}(\tau)$ for 2D binary systems with $\varepsilon=0.8\sigma$ at $\varphi_{c}=0.6324$. This finite-size scaling is different from the typical BRO model with a large step size ($\varepsilon=1.5\sigma$) (see Fig. S2). The obtained critical exponents $\beta$, $\nu_{\parallel}$, $\alpha$, $z$ as functions of $\varepsilon$ are shown in Fig. 2(c, f), where for $\varepsilon>1.0\sigma$ (Regime I), these critical exponents are consistent with the Manna universality. When $0.3\sigma<\varepsilon<1.0\sigma$ (Regime II), we find non-monotonic behaviors in $\nu_{\perp}$, $\alpha$, and $z$ inconsistent with the Manna universality. Further decreasing $\varepsilon$ below $0.3\sigma$ (Regime III), the failure of standard finite size scaling and the uncertainty of the critical point are observed. It should be noted that the exponent $\beta$ does not change significantly in Regimes I and II. The measured $z$, $z^{\prime}$, and $z^{*}$ are consistent, and $\nu_{\perp}=z/\nu_{\parallel}$ also changes with $\varepsilon$ (see Fig. S3). In Fig. S3-6, we also present the exponents measured in the 3D system under two different particle diameter ratios, as well as the monodisperse 4D system. We find that the general behavior of the critical exponent and finite size scaling remain unchanged. This indicates that the anomalous criticality of the BRO model is robust at high density, insensitive to the dimensionality and particle diameter ratio.
@@
We first focus on the Regime I where the step size $\varepsilon$ is larger than the particle size. In this regime, the motion of active particles is not constrained by neighbouring particles. Thus, the mean-square-displacement (MSD) of active particles shows diffusion scaling as shown in Fig. 2(g). Here, the solid and open symbols represent systems above and below the critical point, respectively. The diffusion scaling in sub-critical systems indicates that the absorbing transition is essentially a non-local searching process for a non-overlapping configuration, consistent with Manna (CDP) universality [40, 41].
As $\varepsilon$ decreases, $\varphi_{c}$ increases. Due to the repulsive displacement in the BRO model, the particles are caged by their neighbours. This effect is shown in Fig. 2(h) for systems with $\varepsilon=0.5\sigma$, where one can find an early-time plateau or sub-diffusion regime in MSD for systems near the critical point, a characteristic of the glass-forming liquids [42, 43, 44, 45]. In Fig. 2(i), we plot the measured diffusion coefficient $D$ of active particles and $f_{a}^{\infty}$ as a function of $\varphi$ under different $\varepsilon$. We can clearly distinguish active states with $D$ as low as $10^{-6}\sigma^{2}/\tau_{0}$ for $\varepsilon\lesssim\sigma$. The decoupling between $f_{a}^{\infty}$ and $D$ suggests that in this regime (Regime II), the absorbing phase transition is essentially an absorbing to active-glass transition, where particles can only find a non-overlapping absorbing configuration through local arrangements instead of non-local searching.
The fundamental difference between these two scenarios can be distinguished by field theory. For Manna (CDP) universality, it can be described by a non-conserved field $\rho({\bf{r}},t)$ (active particle density field) coupling with a non-diffusive conserved field $\psi({\bf{r}},t)$ (total density field) [46, 47], namely,
@@
with $\theta=1$ indicating the evolution of $\psi(t)$ field is a result of diffusion of $\rho(t)$ field. Note that $a$, $b$, $w$ are fixed parameters determining the critical point [47]. The last term in Eq.(4) is the multiplicative noise term accounting for absorbing phase transition. In the active-glass state, however, the dynamic of $\psi(t)$ field in Eq.(4) is sub-diffusive, corresponding to fractional time derivative with $0<\theta<1$ [48, 49], while the activity field is still diffusive as activity can spread even when particles are caged. Obviously, the criticality with fractional time dynamics distinguishes itself from Manna universality. The observed continuous variation of critical exponents in this regime might be due to the variation of $\theta$. Nevertheless, numerical solution and renormalization-group analysis of this fractional time equation with multiplicative noise currently remains challenging [48, 49].
@@
When $\varepsilon$ is further reduced below $0.3\sigma$ (Regime III), the critical density of the system would approach that of RCP state. In this regime, caging-breaking events become unlikely and the topology of particle network remains unchanged in the activation process. We refer to this state as the pre-jamming state. Different from the conventional absorbing phase transition, we find $f_{b}(t)$ in this state relaxes according to a continuous power law
@@
In the Griffiths phase, the quenched disorder generates strong dynamic heterogeneity in the relaxation process. This can be visualized by the spatial distribution of normalized activity duration $f_{ad}$, i.e., the fraction of time that a particle stays in an active state relative to the active duration of the whole system, starting from a perturbed absorbing configuration. In Fig. 3(d), we show these distributions for systems with $\varepsilon=0.2\sigma$. One can observe non-uniform network structures of activity, resembling the dynamic heterogeneity of supercooled liquids or the force chain of the jammed state [56, 57, 58]. In Regime II, one can also find relatively weak dynamic heterogeneity (Fig. S8), which would be smoothed out by further increasing $\varepsilon$ to Regime I.
In the Griffiths phase, the finite-size effect of $f_{b}(t)$ and $P_{s}(t)$ near criticality is also reversed, i.e., at the same density near the transition, small systems stay longer in the active state than large systems. This is best shown in Fig. 3(b) and Fig. 4(a), as well as in Fig. S2. This abnormality is nevertheless consistent with the fact that small systems are more likely to be jammed near the jamming point [59, 60]. This suggests that close-packing significantly influences the criticality of absorbing phase transition in this regime. Further support comes from the structure factor $S(q)$ for systems with $\varepsilon\rightarrow 0$, which does not follow the critical $q^{0.45}$ scaling of the Manna model, as expected by the field theory (see Fig. S9).
To further prove that the above phenomena are results of the Griffiths effect, we manually add quenched disorder in systems located at Regime I, by making $\varepsilon$ spatially dependent. More explicitly, the system is divided into cells of size $l$. Each cell $(i,j)$ has its own $\varepsilon_{i,j}$ set as
@@
Figure 4: (a, c) Finite size effects of survival probability of dynamic trajectory $P_{s}(t)$. (d, f) Distributions of normalized activity duration. (a, b) are for the original BRO models near the transition point at $\varepsilon=0.01\sigma$ and $\varphi=0.8380$. (c, d) are for the heterogeneous contact model at $P_{e}=0.788$ and $k=1$ based on the same configuration as in (b).
@@
Finally, we study the critical behaviors of the BRO model in the close packing limit ($\varepsilon\rightarrow 0$). In this limit, the particle configuration can be regarded as static during the relaxation process. From the perspective of field theory, this scenario corresponds to $\theta=0$ in Eq.(4), where the evolution of density field $\psi$ is completely forbidden. Thus, $\psi$ acts as pure quenched disorder in the dynamics of $\rho$ field, essentially changing the CDP universality to the directed percolation (DP) with quenched disorder [53, 52, 50]. The classical contact model is the simplest realization of DP [62, 63]. To test this idea, we simulate a contact model on the networks composed of nearly jammed particles generated by the BRO model at $\varepsilon=0.05\sigma$, where each particle represents a node. We find that the topology of the contact network alone does not affect the critical behaviors of systems. However, if the infection rate of a node is proportional to the activity durations of the corresponding particle (see the Method section for details), the activity distribution of this contact model exhibits dynamic heterogeneity resembling that of the high density BRO model, as shown in Fig. 4(b, d). Correspondingly, the finite size scaling of $P_{s}(t)$ in Fig. 4(c) also exhibits similar anomalous behaviors as the BRO model in the close packing limit in Fig. 4(a). This similarity proves that the critical behaviors of the BRO model in the nearly jammed configuration correspond to DP on hierarchical heterogeneous network. Moreover, the Griffiths effect becomes stronger as the heterogeneity increases in the DP model, i.e., for $k=2,4$. (see Fig. S10). It should be noted that our previous modified BRO model with cell-size quenched-disorder fails to account for the Griffiths effect at the close packing limit, because the correlation at this point spans the system size, leaving the topology of heterogeneous network to play the dominant role.